OpenStudy (anonymous):

Which A. y=5x B. y=6/x C. y=0.7x D. y=x/9

OpenStudy (midhun.madhu1987):

Which???

OpenStudy (anonymous):

Which does not show a direct variation between x and y?

OpenStudy (anonymous):

sorry

OpenStudy (midhun.madhu1987):

Thats okay :)

OpenStudy (anonymous):

i thought it was it but i just wasnt even 10% sure :P

OpenStudy (midhun.madhu1987):

So they are inversely proportional

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

4 more?

OpenStudy (midhun.madhu1987):

:0

OpenStudy (midhun.madhu1987):

sure lol

OpenStudy (anonymous):

Which equations show a direct variation between x and y when solved for y? Choose all answers that are correct. A. y-9=2x+9 B. y=4/7 x C. 4+y=1/3 x+4 D. y=3/x

OpenStudy (midhun.madhu1987):

Since we did the previous question... 2 options can be deleted.. which are they?

OpenStudy (anonymous):

umm... what to options?

OpenStudy (midhun.madhu1987):